OpenStudy (anonymous):

Graph the equation with a diameter that has endpoints at (-3, 4) and (5, -2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):

@zpupster

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

helpp? please! will give metals!!

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

What is the distance between these 2 points?

OpenStudy (anonymous):

idk. it doesnt give me that.

OpenStudy (johnweldon1993):

...That's because you're supposed to find it...but we'll do that later...the general equation of a circle looks like \[\large (x - h)^2 + (y - k)^2 = r^2\] Thecenter of the circle is at point (h,k) lets find that...it would be the midpoint between the 2 endpoints we were given so use the fact that \[\large Midpoint = (\frac{x_! + x_2}{2}),(\frac{y_1 + y_2}{2})\] Plug in your 2 coordiantes into that equation and solve for the 'x' and the 'y' coordinates of the midpoint

OpenStudy (anonymous):

(1,1)?

OpenStudy (johnweldon1993):

Correct...the center (or point (h,k) would be (1,1)) So now...back to our equation of a circle.... \[\large (x - 1)^2 + (y - 1)^2 = r^2\] Now we want to solve for 'r' here....so pick one of your 2 points...

OpenStudy (anonymous):

(-3,4)?

OpenStudy (johnweldon1993):

Okay...that (-3,4) will now be plugged in for 'x' and 'y' in that circle equation \[\large (-3 - 1)^2 + (4 - 1)^2 = r^2\] \[\large (-4)^2 + (3)^2 = r^2\] \[\large 16 + 9 = r^2\] \[\large 25 = r^2\] Make sense how I simplified that? So now...our original circle equation \[\large (x - h)^2 + (y-k)^2 = r^2\] will now look like \[\large (x - 1)^2 + (y - k)^2 = 25\]

OpenStudy (anonymous):

i think i get i, so when i graph it, the center point would be (-1,0)?

OpenStudy (johnweldon1993):

No the center point will be (1,1) like we found

OpenStudy (anonymous):

oh ok.

OpenStudy (anonymous):

so would the equation be (x-1)2+(y-1)2?

OpenStudy (johnweldon1993):

... = 25 Remember the whole equation is \[\large (x - h)^2 +(y - k)^2 = r^2\] so we need to include the founded r^2 = 25 in the equation.. \[\large (x - 1)^2 + (y - 1)^2 = 25\]

OpenStudy (anonymous):

thank you!